OpenStudy (anonymous):

find the value of lim tan(pi/4 + h)-tan(pi/4)/h as h approaches 0 help :(

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\tan(\frac{\pi}{4}+h)-\tan(\frac{\pi}{4})}{h}\] if you know derivative of trig functions this can be quick if not then this will be a little long

OpenStudy (anonymous):

uhm yeah i know the derivative of trig functions

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}=f'(x) \\ \lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h}=f'(x)|_{x=a}\]

myininaya (myininaya):

notice your f(x)=tan(x) and a is?

myininaya (myininaya):

so just find f'(x) then plug in the a value you have there

OpenStudy (calculusfunctions):

First I need you to use the identity\[\tan (A +B)=\frac{ \tan A +\tan B }{ 1-\tan A \tan B}\]to simplify the limit expression.

OpenStudy (anonymous):

guess knowing the derivative of tangent is out of the question here huh?

OpenStudy (calculusfunctions):

Yes it is same as fin ding the derivative and evaluating however they want you find the derivative from first principles here so even though @myininaya is correct she is misguiding you unfortunately.

OpenStudy (anonymous):

uhm okay i don't quite understand

OpenStudy (calculusfunctions):

What part do you not understand??

OpenStudy (anonymous):

i know what the correct answer is i just don't know how to get it

myininaya (myininaya):

I don't believe I'm misguiding anyone. If they know the derivatives of the trig function and the definition of derivative, then I don't quite see how I'm misguiding.

OpenStudy (calculusfunctions):

Do you not understand the identity I told you to use?

OpenStudy (anonymous):

no i don't

myininaya (myininaya):

And as they said they do know the derivative of the trig functions.

OpenStudy (anonymous):

uhm so @myininaya can you just help me?

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\tan(\frac{\pi}{4}+h)-\tan(\frac{\pi}{4})}{h}=(\tan(x))'|_{x=\frac{\pi}{4}}\] this was the way I suggested

OpenStudy (calculusfunctions):

@dolphins121 I'll show you the first step\[\lim_{h \rightarrow 0}\frac{ \tan (\frac{ \pi }{ 4 }+h)-\tan \frac{ \pi }{ 4 } }{ h }\]\[=\lim_{h \rightarrow 0}\frac{ \frac{ \tan \frac{ \pi }{ 4 }+\tan h }{ 1-\tan \frac{ \pi }{ 4 }\tan h }-\tan \frac{ \pi }{ 4 } }{ h }\]Now can you take over from here.

myininaya (myininaya):

find the derivative of tan(x) then evaluate that derivative at x=pi/4

myininaya (myininaya):

and you are done

OpenStudy (anonymous):

wait okay so the derivative of tanx is sec^2x so what do i do from there?

OpenStudy (calculusfunctions):

I kept loosing my internet connection and had to start over again. Plus they want it done from first principles and not take the derivative directly. That's where @myininaya is misguiding you!! I am a teacher who signs in here whenever I have free time so that's why I can say that @myininaya you are misguiding him.

myininaya (myininaya):

I'm not misguiding anyone. I have been a teacher as well. You aren't the only one who teaches.

myininaya (myininaya):

There are multiple ways to answer a question. Please do not get upset if I come up with another way that is different from yours.

OpenStudy (calculusfunctions):

Yes\[\frac{ d }{ dx }\tan x =\sec ^{2}x\]and you can substitute pi/4 in for x but once again you are supposed to arrive at the answer by using the definition (first principles)

myininaya (myininaya):

And telling someone I'm misguiding them sounds like I'm leading them to a wrong answer and makes it sounds like I'm not even giving them a correct way to follow.

myininaya (myininaya):

Says who?!

OpenStudy (anonymous):

I agree with @myininaya . Her method is shorter and more efficient.

OpenStudy (calculusfunctions):

You're missing the point!!! @myininaya I know there are various ways and you are not wrong but he is supposed to use first principles!!! Go ask your own professor!

OpenStudy (calculusfunctions):

Again you people are missing the point but whatever you don't have to believe me!! If this is an assignment that will be marked, you'll find out the hard way. Good Luck!!

OpenStudy (anonymous):

okay but it doesn't say that i need to use first principles so @myininaya can you keep helping me. i sort of understood i just need some guidance

myininaya (myininaya):

Do you know the derivative of tan(x) @dolphins121 ?

OpenStudy (anonymous):

yes its sec^2x

myininaya (myininaya):

now just plug in pi/4

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\tan(\frac{\pi}{4}+h)-\tan(\frac{\pi}{4})}{h}=(\tan(x))'|_{x=\frac{\pi}{4}} \\ =\sec^2(x)|_{x=\frac{\pi}{4}}=\sec^2(\frac{\pi}{4})= [\sec(\frac{\pi}{4})]^2\]

myininaya (myininaya):

do you know cos(pi/4) because sec(pi/4) will the reciprocal of that

OpenStudy (anonymous):

yeah it would be 1 divided by sqrt2/2

myininaya (myininaya):

\[(\frac{1}{\frac{\sqrt{2}}{2}})^2=(\frac{2}{\sqrt{2}})^2\]

myininaya (myininaya):

then you have 2^2/2

myininaya (myininaya):

I will let you finish that

OpenStudy (anonymous):

okay so then that simplifies to 2

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{\tan(\frac{\pi}{4}+h)-\tan(\frac{\pi}{4})}{h}=(\tan(x))'|_{x=\frac{\pi}{4}} \text{ by definition of derivative } \\ =\sec^2(x)|_{x=\frac{\pi}{4}}=\sec^2(\frac{\pi}{4})= [\sec(\frac{\pi}{4})]^2 \] \[=(\frac{2}{\sqrt{2}})^2=\frac{2^2}{2}=\frac{4}{2}=2\] we can add a little note up there so no points will taken

OpenStudy (anonymous):

okay but wouldn't the whole thing be zero because you have sec^2(pi/4+h)-sec^2(pi/4)/h and when you work everything out you'll be subtracting 2-2? and since h is approaching zero the denominator is zero anyways?

myininaya (myininaya):

are you familiar with this: \[\lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h}=f'(a)\]

myininaya (myininaya):

this is the definition of derivative

OpenStudy (anonymous):

yes

myininaya (myininaya):

we had f(x)=tan(x) and a=pi/4

myininaya (myininaya):

\[f(\frac{\pi}{4}+h)=\tan(\frac{\pi}{4}+h) \\ f(\frac{\pi}{4})=\tan(\frac{\pi}{4}) \\\]

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h}=f'(a) \\ \lim_{h \rightarrow 0} \frac{\tan(\frac{\pi}{4}+h)-\tan(\frac{\pi}{4})}{h}=(\tan(a))'|_{a=\frac{\pi}{4}}=\sec^2(a)|_{x=\frac{\pi}{4}}=\sec^2(\frac{\pi}{4})=2\] do you that our limit expression is in the same form where f(x)=tan(x) and a=pi/4

OpenStudy (anonymous):

uhm yeah i get that. but don't we plug in zero for h? that's what's confusing me

myininaya (myininaya):

we don't have any h's anymore when we get to the derivative of tan(x) part

OpenStudy (anonymous):

oh. okay so can you just show me how it would look to see if i understand

myininaya (myininaya):

But I already showed you \[\lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h}=[f(x)]'|_{x=a}\ \text{ we have } f(x)=\tan(x) \text{ and } a=\frac{\pi}{4} \\ \lim_{h \rightarrow 0}\frac{\tan(\frac{\pi}{4}-h)-\tan(\frac{\pi}{4})}{h}=[\tan(x)]'|_{x=\frac{\pi}{4}} \text{ note: no h's here }\]

myininaya (myininaya):

say we have this limit expression \[\lim_{h \rightarrow 0}\frac{\sqrt{4+h}-\sqrt{4}}{h}=[\sqrt{x}]'_{x=4}=\frac{1}{2\sqrt{x}}|_{x=4}=\frac{1}{2 \sqrt{4}}=\frac{1}{2(2)}=\frac{1}{4}\]

myininaya (myininaya):

or you could rationalize the numerator and still wind up with 1/4 \[\lim_{h \rightarrow 0}\frac{4+h-4}{h(\sqrt{4+h}+\sqrt{4})}=\lim_{h \rightarrow 0}\frac{1}{\sqrt{4+h}+\sqrt{4}}=\frac{1}{\sqrt{4+0}+\sqrt{4}}=\frac{1}{2+2}=\frac{1}{4}\]

myininaya (myininaya):

still makes no sense?

OpenStudy (anonymous):

oh okay. i get it. thanks

myininaya (myininaya):

here is another example just in case \[\lim_{h \rightarrow 0}\frac{\cos(\frac{\pi}{4}+h)-\cos(\frac{\pi}{4})}{h}=(\cos(x))'_{x=\frac{\pi}{4}}=-\sin(x)|_{x=\frac{\pi}{4}}=-\sin(\frac{\pi}{4})=- \frac{\sqrt{2}}{2}\]

myininaya (myininaya):

so whenever you have \[\lim_{h \rightarrow 0}\frac{f(a+h)-f(a)}{h} \text{ you can skip to } (f(x))'_{x=a}\]

OpenStudy (anonymous):

oh okay. well thanks

OpenStudy (zarkon):

Well that was fun ;)